OpenStudy (anonymous):

Given the linear equation 4y = 8x – 3, find the y-intercept.

OpenStudy (anonymous):

At y intercept x=0

OpenStudy (anonymous):

Substitute x=0 and get ur answer!!

OpenStudy (anonymous):

not an option im thinking its -3/4 or 3/4

OpenStudy (anonymous):

why two answers??

OpenStudy (anonymous):

its not quadratic

OpenStudy (anonymous):

because im not sure if it would be negative or possitive

OpenStudy (anonymous):

just tell me what u get if u substitute x=0??

OpenStudy (anonymous):

4y=8(0)-3 y=-3/4

OpenStudy (anonymous):

Let x=0

OpenStudy (anonymous):

there u go....its negative.......

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Congrats........and always get the exact same value u get after substitution!!

OpenStudy (anonymous):

ur welcome!!